20170131 » History » Version 21
Nebrensky, Henry, 09 February 2017 11:27
Next meeting
1 | 1 | Boyd, Steven | h1. Ops/Comm Scheduling meeting |
---|---|---|---|
2 | |||
3 | 2 | Boyd, Steven | Location : G06 in R66 at 2pm |
4 | 1 | Boyd, Steven | |
5 | Phone access : ReadyTalk number |
||
6 | US 866-740-1260 |
||
7 | UK 0800-496-0576 |
||
8 | Swiss 0800-70-53-51 |
||
9 | Italy 800182592 |
||
10 | Access Code: 2155402 |
||
11 | |||
12 | * SS Magnet status |
||
13 | * Hydrogen status |
||
14 | * Issues arising from previous running |
||
15 | 4 | Boyd, Steven | ** [[Hall_Task_List_Dec16Feb17|Shutdown work in Jan and Feb]]: "present version":https://micewww.pp.rl.ac.uk/attachments/download/8158 |
16 | 1 | Boyd, Steven | * Operations : |
17 | 6 | Boyd, Steven | ** C&M System Check report (Ajit/Ed) |
18 | 1 | Boyd, Steven | ** Establishing a timetable for the runup : when should systems (including magnets) be ready for running? |
19 | *** Beam schedule |
||
20 | **** Run up on Jan 30 - Feb 12. |
||
21 | **** Start of user cycle : Feb 14 |
||
22 | **** Start of data-taking : Feb 15 |
||
23 | 14 | Boyd, Steven | *** Reserve shifter support from Feb 10 - Feb 14 |
24 | 1 | Boyd, Steven | *** MOM formally starts on Feb 9 (but should be present for the runup activities but will not be on 24-hour call) |
25 | *** Possible runup plan (framework for discussion) |
||
26 | 7 | Boyd, Steven | **** Monday 6th : Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running |
27 | 15 | Boyd, Steven | **** Tuesday 7th (subject to meeting on Monday) : Bring up magnets to running configuration, take some pulser data and take magnets down |
28 | 5 | Boyd, Steven | **** Saturday 11th : Expert led run without solenoids + Beam tune |
29 | **** Monday 13th : Triage actions arising from expert led run |
||
30 | 8 | Boyd, Steven | **** Tuesday 14th : Bring up magnets in the evening and monitor overnight by reserve shift. |
31 | 16 | Boyd, Steven | **** Wednesday 15th : Data-taking start at 08:00 |
32 | 1 | Boyd, Steven | |
33 | ** Shift signup for first half of Feb is viable now (but not full) |
||
34 | |||
35 | 10 | Boyd, Steven | | *Shift Block ID* | # signed up | # shifters still needed | |
36 | 1 | Boyd, Steven | | 1 | 6 | 0 | |
37 | 11 | Boyd, Steven | | 2 | 4 | 2 | *Need one overnight shift per night (5 nights) for monitoring* | |
38 | | 3 | 4 | 2 | *Need one overnight shift per night (5 nights) for monitoring* | |
||
39 | 1 | Boyd, Steven | | 4 | 4 | 1 | |
40 | |||
41 | 12 | Boyd, Steven | *Magnet monitors do not need to be trained shifters* |
42 | |||
43 | 13 | Boyd, Steven | ** BLOCs are needed : Feb 20 - 24 and Feb 27 - Mar 1 |
44 | 1 | Boyd, Steven | |
45 | * AOB: |
||
46 | |||
47 | * Last meeting : [[20170124|Tuesday Jan 24 2017]] |
||
48 | 21 | Nebrensky, Henry | * Readiness meeting : [[20170206|Monday Feb 6 2017]] |
49 | * Next meeting : [[20170207|Tuesday Feb 7 2017]] |
||
50 | 17 | Nebrensky, Henry | |
51 | h2. Minutes |
||
52 | |||
53 | * CC Magnet status (J Boehm) |
||
54 | All 3 CC magnets are now in flip mode; cold and full of helium. Everything is presently being run into shorts for the CC IOC testing. |
||
55 | Handover from Pierrick is in progress. A few GUIs have been changed. |
||
56 | Are seeing instability in SSU power supply (into shorts): last week suddenly ramped from 150 to 215 A; lost control of GUI; ramped down manually. |
||
57 | Similar last night, with control cables disconnected from the AMI controller - again unstable and major changes in current on its own, implies a hardware issue. |
||
58 | Now running with a different controller - seeing issues still. May possibly be associated with a purely resistive load. |
||
59 | |||
60 | * Hydrogen status (M Tucker) |
||
61 | Cooled down in R9 with Neon, made 2 to 2.5 litres of liquid, less than 1 K difference between condenser and absorber. |
||
62 | Expect about 1 l every 6 hours, so over 5 days to fill with LH2 needing a 24hour rota. |
||
63 | |||
64 | * Issues arising from previous running (DC) |
||
65 | Need a comprehensive list of encountered issues, else can't assume they will be addressed before running resumes. We are now well past halfway through the shutdown. |
||
66 | * Shutdown work in Jan and Feb ( "official version":https://micewww.pp.rl.ac.uk/attachments/download/8158 ) |
||
67 | Issues/Progress on workplan items (using a stale list): |
||
68 | |||
69 | ** PT100s - done |
||
70 | ** ODH sensor service - done |
||
71 | ** POMPOMs sensor test - done |
||
72 | ** Neutron Monitor: ISIS can't do it; will return to MICE for us to make arrangements. Continue with temporary monitor this cycle. |
||
73 | ** Proton Absorbers - remedial work done |
||
74 | ** "New" chiller plumbing - done |
||
75 | ** #1878 Trench heat exchanger to counter-flow - done, possibly some small leaks |
||
76 | |||
77 | ** Diffuser iris 4: needs next SSU run: Tuesday 7th |
||
78 | Compressed air: Activair coming to check the outside compressor to see why it didn't come on. |
||
79 | |||
80 | ** #1399 MICEnet lockout - tomorrow morning |
||
81 | |||
82 | ** #1888 Diffuser air - reported done |
||
83 | ** Airflow monitoring changes - done |
||
84 | ** PPS Functional Test - done |
||
85 | ** Earth leakage tests - done |
||
86 | ** New @target1ctl@ - issues with O/S and libraries... |
||
87 | ** R78 - replacement flanges in progress |
||
88 | |||
89 | 20 | Nebrensky, Henry | * The Decay Solenoid is cold. Only known issue is the episodes of fridge instability seen last run. |
90 | Presently the cooling lead flow is on, to check the stability of the fridge; otherwise ready to power. |
||
91 | |||
92 | 17 | Nebrensky, Henry | * C&M System Check report (Ajit/Ed) |
93 | CC Magnet UIs/IOCs tested last week - changes have been made on various machines, need to be aggregated. |
||
94 | Target SM tested today. |
||
95 | Need to test Run Control: Thursday? |
||
96 | Need to review the tags listed in the CDB |
||
97 | The NIKHEF Hall probes have been separated out into a different IOC. |
||
98 | Most of the hardware stuff is done. |
||
99 | |||
100 | * Run-up plan |
||
101 | ** Roof water/chiller needs a load test with the air-blast active - tomorrow or Thursday subject to Hall work and state of pipes |
||
102 | ** Monday 6th: Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running; remove shorts from link boxes, ATW and other prep. |
||
103 | ** Tuesday 7th: Bring up CC magnets to running configuration, take some pulser data and take magnets down; _check Diffuser operation_ |
||
104 | ** Saturday 11th - Half-day of ISIS beamtime: expert led run without CC Magnets + Beambump tune (Hayley, NB Ed has to leave mid-afternoon) |
||
105 | ** Monday 13th: Triage actions arising from expert led run; BLOC training on everything that isn't target |
||
106 | ** Tuesday 14th: Bring up magnets in the evening and soak-test overnight (reserve shifters). |
||
107 | ** Wednesday 15th: Data-taking start at 08:00 |
||
108 | |||
109 | * Shift signup: |
||
110 | 18 | Nebrensky, Henry | With the LH2 install going ahead there will be no running during second half of 2016/05 and thus no available shift allocation - MICE that have not yet made shift quota take note. |
111 | 17 | Nebrensky, Henry | |
112 | * AOB: |
||
113 | Linde Gas Analyser webcam - where has it gone? It is being PAT tested. |
||
114 | |||
115 | h3. Actions |
||
116 | |||
117 | * JBoehm to get SS operating info from PHanlet (ongoing) |
||
118 | * JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI" |
||
119 | * DC to collate list of issues arising from previous running |
||
120 | * Status of @miceraid5@ disk cleanup (stands) |
||
121 | * -AKurup to make and circulate plan for NIKHEF probe IOC- |
||
122 | * -SBoyd - ISIS machine-physics beam time- |
||
123 | * DC to ensure Health Physics have checked temporary neutron monitor (stands) |
||
124 | * SBoyd - ensure MICE-wide discussion of MOM C&M panel |
||
125 | 19 | Nebrensky, Henry | * Henry to get gaps in BLOC rota plugged |